OpenStudy (anonymous):

Linear Approximation of ln3?

OpenStudy (anonymous):

I know that e^x will be the closest thing, but I am not sure how to work that into the problem.

OpenStudy (amistre64):

a linear approx is the line equation defined by: y = f'(a)(x-a) + f(a) this simply defines the slope of the line at a (the derivative), using the point a,f(a) as its anchor

OpenStudy (anonymous):

yes, I see, but what point?

OpenStudy (amistre64):

3,ln3 which may defeat the purpose since it is ln3 you are wanting to find :/

OpenStudy (amistre64):

im thinking the taylor polynomial would be the next best thing

OpenStudy (amistre64):

if we could make a polynomial that best fits this curve, we would want it to "move" in the same way as it .... movement is a rate of change, or derivative.

OpenStudy (amistre64):

spose \[f(x)=a_0+a_1x+a_2x^2+a_3x^3+...\] \[f'(x)=a_1+2a_2x+3a_3x^2+...\] \[f''(x)=2a_2+3.2a_3x+...\] all we would have to do is find some way to define the constants .... and we would have a polynomial f(x) that fits the curve

OpenStudy (amistre64):

one strategy we could do is to zero out all the x parts, that would leave us with the constants at the start of each derivative. For x=3, we would want to modify each x to get (x-3) ... in order to zero it out

OpenStudy (amistre64):

\[f(x)=a_0+a_1(x-3)+a_2(x-3)^2+a_3(x-3)^3+...\] \[f'(x)=a_1+2a_2(x-3)+3a_3(x-3)^2+...\] \[f''(x)=2a_2+3.2a_3(x-3)+...\] now when x=3, we get \[f(3)=a_0\] \[f'(3)=a_1\] \[f''(3)=2a_2\] \[...\] and we can solve for the coefficients

OpenStudy (amistre64):

any of this making sense?